OpenStudy (anonymous):

Don't even know where to start D: Solve 3^x – 4 = 7^x + 9 Please and thank you!

OpenStudy (anonymous):

group like terms

OpenStudy (anonymous):

What would that look like?

OpenStudy (anonymous):

grouping like terms isn't the proper wording

OpenStudy (anonymous):

wat shuld we say \[3^x-7^x=13\]

OpenStudy (anonymous):

okay, then? Im sorry im not getting it at all.

OpenStudy (anonymous):

is this your question \[3^x-4=7^x+9\]